OpenStudy (anonymous):

Please help! Greg has only nickels and dimes in his money box. He knows that he has more than $10 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y?

OpenStudy (anonymous):

Draw a dashed line to represent the graph of 5x + 10y = 1000, and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 1000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 1000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 1000, and shade the portion below the line for positive values of x and y.